# How Global Calibration Strengthens Multiaccuracy

### v1

@@ 5.4 Density of $\mu_{\mathsf{TTV}}$ vs $\mu_{\mathsf{Max}}$
We conclude this section with an example illustrating the density difference between $\mu_{\mathsf{TTV}}$ and $\mu_{\mathsf{Max}}$. Consider a domain $\mathcal{X}$ and a probability distribution $\mathcal{D}$ on $\mathcal{X}$. We define a predictor $p:\mathcal{X}\rightarrow[0,1]$ that predicts only two values on $\mathcal{X}$: $p(x)=1/2$ and $p(x)=\eta$ for some small value $20<\eta<1/$ 2. Let $S_{1/2}=\{x:p(x)=1/2\}$ and $S_{\eta}=\{x:p(x)=\eta\}$; their sizes are such that $\Pr_{\mathcal{D}_{\mathcal{X}}}[\mathbf{x}\in S_{1/2}]=2\delta$ and $\Pr_{\mathcal{D}_{\mathcal{X}}}[\mathbf{x}\in S_{1/2}]=1-2\delta$, where $\delta>0$ and $\eta\ll\delta$. Moreover, $p$ is perfectly calibrated for $\mathcal{D}^{g}$, where $g$ is the boolean input function. Given these $p,g$, and $\mathcal{D}$, we compare the densities of the measures $\mu_{\mathsf{TTV}}$ and $\mu_{\mathsf{Max}}$. Figure 3 summarizes the construction (drawn for example values $\eta=0.05,\delta=0.2$).
@@
  • $\mathbf{\mu_{\mathsf{TTV}}}:$ In $S_{1/2}$, $p=1-p=1/2$, and so $\mu_{\mathsf{TTV}}$ assigns weight $1/2$ to all $\mathbf{x}\in S_{1/2}$. By the calibration guarantee on $p$, within $S_{1/2}$ half the points have $g$-value equal to 0 and half the points have $g$-value equal to 1. In $S_{\eta}$, $\mu_{\mathsf{TTV}}$ gives weight $p=\eta$ to the points where $g(x)=0$ and weight $1-p=1-\eta$ to the points where $g(x)=1$. By the calibration guarantee on $p$, within $S_{\eta}$ a $(1-\eta)$-fraction of points have $g$-value equal to 0, and a $\eta$-fraction of points have $g$-value equal to 1. Therefore,
@@
  • $\mathbf{\mu_{\mathsf{Max}}}:$ In $S_{1/2}$, $p=1-p$, so $\mu_{\mathsf{Max}}$ assigns weight 1 to all $\mathbf{x}\in S_{1/2}$. In $S_{\eta}$, $\mu_{\mathsf{Max}}$ gives weight $\eta(1-\eta)$ to the points where $g(x)=0$ weight $1$ to the points where $g(x)=1$. Again using the calibration guarantee on $p$ within each of $S_{1/2}$ and $S_{\eta}$ it follows that
@@ 6 Restricted Weak Agnostic Learning and Multiacccuracy
Theorem 4.2 proves a reduction to multiaccuracy from $(\alpha,\beta)$-weak agnostic learning for $\alpha>1/2$. This motivates the study of weak learning in the setting where $\alpha$ is bounded away from 00. In this section, we investigate the notion of restricted weak agnostic learning, i.e., $(\alpha,\beta)$-weak agnostic learning when $\alpha$ is bounded away from 00, and study what learning and fairness primitives can be derived from such restricted weak agnostic learners.
@@
The agnostic boosting results of [13, 30] tell us that increasing $\beta$ does not necessarily result in harder learning problems. They show that an $(\alpha,\beta)$ weak learner implies a $(\gamma,\gamma-\alpha)$ weak learner for all $\gamma>\alpha$. This result is typically interpreted to say that if weak agnostic learning is possible for $\alpha$ approaching 00, then strong agnostic learning is possible.
@@
Our construction of classes that are hard to learn assumes the existence of pseudorandom families of functions, and on the construction of binary error-correcting codes that can be (list)decoded from error rates approaching $1/2$ [22]. Decoding from error rate $1/2$ is information-theoretically impossible, since a random string would be at distance $1/2$ from a codeword. The gap between distance $1/2$ and $1/2-\varepsilon$ can be viewed as a gap of 00 versus $2\varepsilon$ in the correlation between the received word and codeword. Our result uses pseudorandom functions to translate this to a computational gap in the learnability of a suitably defined concept class with correlation $\alpha$ versus $\alpha+\varepsilon$. Our construction is inspired by a construction of Feldman [12] which he uses to show that membership queries add power to agnostic learning.
@@
There exists a family of pseudorandom functions, denoted =$\mathcal{F}=\{f_{r}:\{0,1\}^{n}\to\{0,1\}~{}|~{}{r\in\{0,1\}^{n}}\}$, such that
  • $f_{r}$ is easy to compute given $r$ as input.
  • There exists a function $T:\mathbb{N}\rightarrow\mathbb{N}$, such that for $r\leftarrow\{0,1\}^{n}$ chosen uniformly at random, any (potentially randomized) algorithm that runs in time $T(n)$ and has query access to $f_{r}:\{0,1\}^{n}\to\{0,1\}$ cannot distinguish it from a truly random function with advantage greater than $1/T(n)$ over the uniform distribution on $x\in\{0,1\}^{n}$.
@@
We then define the {±$\{\pm 1\}$ version of $g_{r}$ as $c_{r}(\mathbf{x})=2g_{r}(\mathbf{x})-1$ and let $\mathcal{C}=\{c_{r}~{}|~{}{r\in\{0,1\}^{n}}\}$.
@@
Using the standard translation from correlation to Hamming distance for variables in {±$\{\pm 1\}$,
@@
$=\operatorname*{\mathbf{E}}[(2p_{r}(\mathbf{x})-1)(2\mathbf{y}-1)|\mathbf{x}\in E]$
$=1-2\Pr[\mathbf{y}\neq p_{r}(\mathbf{x})|x\in E].$
Since the LHS is at least $\varepsilon/(1-\alpha)$, and $p_{r}(x)$ equals $\mathrm{RSH}(r)$ on the set $E$, we conclude that
@@
$=\sum_{x\in[m]}\Pr[\mathbf{y}\neq\mathrm{RSH}(r)_{x}|\mathbf{x}^{*}=x]$
@@
Hence by a Chernoff bound, with probability $1-\exp(-\varepsilon^{2}\alpha^{2}m)$ we have $d(\mathbf{w},\mathrm{RSH}(r))\leq(1-\varepsilon)m/2$. So running the list-decoder from random samples returns a list containing the seed $r$. We can enumerate over all candidates $r^{\prime}$ in the list, and check whether $\mathrm{cor}(\mathbf{y},c_{r^{\prime}})\geq\alpha$ using random samples, using the fact that $f_{r}$ is easy to compute given $r$. Taking $r^{\prime}=r$ will satisfy give $\alpha+\varepsilon$ correlation, but there might be better hypotheses. We output the best hypothesis that we find, which will have correlation at least $\alpha$ with probability $0.9$. [^16] We lose an $\varepsilon$ in accounting for sampling error in estimating the true correlation. We can amplify the probability of success through repetition.
@@
where we use $c_{r}(x)=2f_{r}(x)=1$ for $x\in F$. Thus, in either case we can predict the pseudorandom function $f_{r}$ with non-trivial advantage with non-negligible probability. This concludes the proof of the result. ∎
@@ 7 Conclusion and Future Work
Another natural question is whether more general reductions to multiaccuracy can yield a weak agnostic learner. The result in Theorem 4.1 only rules out learners obtained by post-processing a predictor satisfying certain multigroup guarantees. A stronger model would be to consider an algorithm that has oracle access to a learner that for multiaccurate predictors, which it can call multiple times. In Appendix A, we show how a more general post-processing involving points $\mathbf{x}\in\mathcal{X}$ in addition to just the value $p(\mathbf{x})$ can in some cases yield agnostic learners.
@@ Acknowledgments
We thank Vitaly Feldman, Venkatesan Guruswami, and Salil Vadhan for helpful discussions.
@@ Appendix A Projecting Multiaccurate Predictors onto the Span of $\mathcal{C}$
Intuitively, the reason why $\mathcal{C}$-multiaccuracy does not directly imply weak agnostic learning for the class $\mathcal{C}$ is that multiaccuracy only implies a certain type of closeness between the predictor $p$ and the labels $\mathbf{y}$ over the span of the concept class $\mathcal{C}$. Let $p^{*}$ denote the Bayes optimal predictor, i.e., $p^{*}(\mathbf{x})=\operatorname*{\mathbf{E}}[\mathbf{y}|\mathbf{x}]$, and suppose for now that $p$ satisfies $(\mathcal{C},0)$-multiaccuracy for $\mathcal{D}$. Let $\tilde{p}^{*}=2p^{*}-1$ and $\tilde{p}=2p-1$ denote the $[-1,1]$ versions of the predictors $p^{*}$ and $p$. Let $\mathsf{span}(\mathcal{C})$ denote the vector space spanned by $\mathcal{C}$. We can write,
@@
where $\tilde{p}_{C},\tilde{p}^{*}_{C}$ are the components of $\tilde{p}$ and $\tilde{p}^{*}$ in the $\mathsf{span}(\mathcal{C})$ respectively, and $\tilde{p}_{\perp},\tilde{p}^{*}_{\perp}$ are the components of $\tilde{p},\tilde{p}^{*}$ orthogonal to $\mathsf{span}(\mathcal{C})$. We recall that the inner product $\langle f,g\rangle$ for functions $f,g:\mathcal{X}\to\mathbb{R}$ is defined by by $\operatorname*{\mathbf{E}}_{\mathcal{D}_{X}}[f(\mathbf{x})g(\mathbf{x})]$. Since $p$ is $(\mathcal{C},0)$-multiaccurate for $\mathcal{D}$, we have that,
$\langle c,\tilde{p}-\tilde{p}^{*}\rangle=2\operatorname*{\mathbf{E}}[c(p(\mathbf{x})-p^{*}(\mathbf{x}))]=0$
@@
As $p_{\perp},p^{*}_{\perp}$ are both orthogonal to $\mathsf{span}(\mathcal{C})$, and $\tilde{p}_{C}-\tilde{p}^{*}_{C}\in\mathsf{span}(\mathcal{C})$, we must have $\tilde{p}_{C}=\tilde{p}^{*}_{C}$. What this suggests is that even though technically we can have
@@
$\mathrm{cor}(\mathbf{y},\tilde{p}_{C})=\langle\tilde{p}^{*}\mathbf{x},\tilde{p}_{C}\rangle$ $=\langle\tilde{p}^{*}_{C}+\tilde{p}^{*}_{\perp},\tilde{p}_{C}\rangle$
@@
Above in the last line we used that $\tilde{p}^{*}_{\perp}$ is orthogonal to $\mathsf{span}(\mathcal{C})$ and that $\tilde{p}_{C}\in\mathsf{span}(\mathcal{C})$, as well as the observation that $\tilde{p}_{C}=\tilde{p}^{*}_{C}$. Thus, a projection of $\tilde{p}$ onto $\mathsf{span}(\mathcal{C})$ could potentially lead to a weak agnostic learner. However, we need to take some care as (a) $\tilde{p}_{C}$ need not have range $[-1,1]$ (or be bounded even), and (b) if we don’t have $(\mathcal{C},0)$-multiaccuracy, but only $(\mathcal{C},\tau)$, then depending on the representation of $\tilde{p}_{C}$ in terms of $c\in C$, the errors could accumulate making the above analysis trivial. Observe that this constitutes a much richer form of post-processing that considered in Section 4.
@@
For simplicity, let $\tilde{p}=2p-1$ denote the $[-1,1]$-version of $p$. Then,
@@
Above in the second last inequality, we used the fact that $p$ is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$ and that $c_{s}\in\mathcal{C}$ for all $s$, and in the last inequality we use that clipping $q$ to get $h\in[-1,1]$ only reduces the squared error with respect to $12\mathbf{y}-$ 1 and that $2\mathbf{y}-1\in\{-1,1\}$. The conclusion then follows by rearranging. ∎
@@
for $\gamma\geq 4\tau\sum_{s}|\lambda_{s}|+\beta$. Then, $h=\operatorname{clip}(q)$, which clips $q$ to the range $[-1,1]$, satisfies $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},h)\geq\beta$.
@@
In the last inequality, we used $\mathrm{cor}(p,c)\geq\alpha$ and $\|c\|\leq 1$, as $c$ has range $[-1,1]$.

### v2

@@ 5.4 Density of $\mu_{\mathsf{TTV}}$ vs $\mu_{\mathsf{Max}}$
We conclude this section with an example illustrating the density difference between $\mu_{\mathsf{TTV}}$ and $\mu_{\mathsf{Max}}$. Consider a domain $\mathcal{X}$ and a probability distribution $\mathcal{D}$ on $\mathcal{X}$. We define a predictor $p:\mathcal{X}\rightarrow[0,1]$ that predicts only two values on $\mathcal{X}$: $p(x)=1/2$ and $p(x)=\eta$ for some small value $0<\eta<1/2$. Let $S_{1/2}=\{x:p(x)=1/2\}$ and $S_{\eta}=\{x:p(x)=\eta\}$; their sizes are such that $\Pr_{\mathcal{D}_{\mathcal{X}}}[\mathbf{x}\in S_{1/2}]=2\delta$ and $\Pr_{\mathcal{D}_{\mathcal{X}}}[\mathbf{x}\in S_{1/2}]=1-2\delta$, where $\delta>0$ and $\eta\ll\delta$. Moreover, $p$ is perfectly calibrated for $\mathcal{D}^{g}$, where $g$ is the boolean input function. Given these $p,g$, and $\mathcal{D}$, we compare the densities of the measures $\mu_{\mathsf{TTV}}$ and $\mu_{\mathsf{Max}}$. Figure 3 summarizes the construction (drawn for example values $\eta=0.05,\delta=0.2$).
@@
  • $\mathbf{\mu_{\mathsf{TTV}}}:$ In $S_{1/2}$, p=1p=1/2p=1-p=1/2, and so $\mu_{\mathsf{TTV}}$ assigns weight $1/2$ to all $\mathbf{x}\in S_{1/2}$. By the calibration guarantee on $p$, within $S_{1/2}$ half the points have $g$-value equal to 0 and half the points have $g$-value equal to 1. In $S_{\eta}$, $\mu_{\mathsf{TTV}}$ gives weight $p=\eta$ to the points where $g(x)=0$ and weight $1-p=1-\eta$ to the points where $g(x)=1$. By the calibration guarantee on $p$, within $S_{\eta}$ a $(1-\eta)$-fraction of points have $g$-value equal to 0, and a $\eta$-fraction of points have $g$-value equal to 1. Therefore,
@@
  • $\mathbf{\mu_{\mathsf{Max}}}:$ In $S_{1/2}$, p=1pp=1-p, so $\mu_{\mathsf{Max}}$ assigns weight 1 to all $\mathbf{x}\in S_{1/2}$. In $S_{\eta}$, $\mu_{\mathsf{Max}}$ gives weight $\eta(1-\eta)$ to the points where $g(x)=0$ weight $1$ to the points where $g(x)=1$. Again using the calibration guarantee on $p$ within each of $S_{1/2}$ and $S_{\eta}$ it follows that
@@ 6 Restricted Weak Agnostic Learning and Multiacccuracy
Theorem 4.4 proves a reduction to multiaccuracy from $(\alpha,\beta)$-weak agnostic learning for $\alpha>1/2$. This motivates the study of weak learning in the setting where $\alpha$ is bounded away from 0. In this section, we investigate the notion of restricted weak agnostic learning, i.e., $(\alpha,\beta)$-weak agnostic learning when $\alpha$ is bounded away from 0, and study what learning and fairness primitives can be derived from such restricted weak agnostic learners.
@@
The agnostic boosting results of [13, 30] tell us that increasing $\beta$ does not necessarily result in harder learning problems. They show that an $(\alpha,\beta)$ weak learner implies a $(\gamma,\gamma-\alpha)$ weak learner for all $\gamma>\alpha$. This result is typically interpreted to say that if weak agnostic learning is possible for $\alpha$ approaching 0, then strong agnostic learning is possible.
@@
Our construction of classes that are hard to learn assumes the existence of pseudorandom families of functions, and on the construction of binary error-correcting codes that can be (list)decoded from error rates approaching $1/2$ [22]. Decoding from error rate $1/2$ is information-theoretically impossible, since a random string would be at distance $1/2$ from a codeword. The gap between distance $1/2$ and $1/2-\varepsilon$ can be viewed as a gap of 0 versus $2\varepsilon$ in the correlation between the received word and codeword. Our result uses pseudorandom functions to translate this to a computational gap in the learnability of a suitably defined concept class with correlation $\alpha$ versus $\alpha+\varepsilon$. Our construction is inspired by a construction of Feldman [12] which he uses to show that membership queries add power to agnostic learning.
@@
There exists a family of pseudorandom functions, denoted $\mathcal{F}=\{f_{r}:\{0,1\}^{n}\to\{0,1\}~|~{r\in\{0,1\}^{n}}\}$, such that
  • $f_{r}$ can be computed in time $\operatorname{poly}(n)$ given $r$ as input.
  • For $r\leftarrow\{0,1\}^{n}$ chosen uniformly at random, any (potentially randomized) algorithm that runs in time $T(n)$ and has query access to $f_{r}:\{0,1\}^{n}\to\{0,1\}$ cannot distinguish it from a truly random function with advantage greater than $1/T(n)$ over the uniform distribution on $x\in\{0,1\}^{n}$.
@@
We then define the $\{\pm 1\}$ version of $g_{r}$ as $c_{r}(\mathbf{x})=2g_{r}(\mathbf{x})-1$ and let $\mathcal{C}=\{c_{r}~|~{r\in\{0,1\}^{n}}\}$.
@@
Using the standard translation from correlation to Hamming distance for variables in $\{\pm 1\}$,
@@
$=\operatorname*{\mathbf{E}}[(2g_{r}(\mathbf{x})-1)(2\mathbf{y}-1)|\mathbf{x}\in E]$
$=1-2\Pr[\mathbf{y}\neq g_{r}(\mathbf{x})|\mathbf{x}\in E].$
Since the LHS is at least $\varepsilon/(1-\alpha)$, and $g_{r}(x)$ equals $\mathrm{RSH}(r)_{x}$ on the set $E$, we conclude that
@@
$=\sum_{x\in[m]}\Pr[\mathbf{y}\neq\mathrm{RSH}(r)_{x}|\mathbf{x}=x]$
@@
Hence by a Chernoff bound, with probability $1-\exp(-\varepsilon^{2}\alpha^{2}m)$ we have $d(\mathbf{w},\mathrm{RSH}(r))\leq(1-\varepsilon)m/2$. So running the list-decoder from random samples returns a list containing the seed $r$. We can enumerate over all candidates $r^{\prime}$ in the list, and check whether $\mathrm{cor}(\mathbf{y},c_{r^{\prime}})\geq\alpha$ using random samples, using the fact that $f_{r}$ is easy to compute given $r$. Taking $r^{\prime}=r$ will give $\alpha+\varepsilon$ correlation, but there might be better hypotheses. We output the best hypothesis that we find, which will have correlation at least $\alpha$ with probability 0.90.9.161616We lose an $\varepsilon$ in accounting for sampling error in estimating the true correlation. We can amplify the probability of success through repetition.
@@
where we use $c_{r}(x)=2f_{r}(x)-1$ for $x\in F$. Thus, in either case we can predict the pseudorandom function $f_{r}$ with non-trivial advantage with non-negligible probability. This concludes the proof of the result. ∎
@@ 7 Conclusion and Future Work
Another natural question is whether more general reductions to multiaccuracy can yield a weak agnostic learner. The result in Theorem 4.1 only rules out learners obtained by post-processing a predictor satisfying certain multigroup guarantees. A stronger model would be to consider an algorithm that has oracle access to a learner that for multiaccurate predictors, which it can call multiple times. In Appendix A, we show how a more general post-processing involving points $x\in\mathcal{X}$ in addition to just the value $p(x)$ can in some cases yield agnostic learners.
@@ Acknowledgments
We thank Igor Carboni Oliveira, Vitaly Feldman, Venkatesan Guruswami, Rahul Santhanam, and Salil Vadhan for helpful discussions and feedback.
@@ Appendix A Projecting Multiaccurate Predictors onto the Span of $\mathcal{C}$
Intuitively, the reason why $\mathcal{C}$-multiaccuracy does not directly imply weak agnostic learning for the class $\mathcal{C}$ is that multiaccuracy only implies a certain type of closeness between the predictor $p$ and the labels $\mathbf{y}$ over the span of the concept class $\mathcal{C}$. Let $p^{*}$ denote the Bayes optimal predictor, i.e., $p^{*}(\mathbf{x})=\operatorname*{\mathbf{E}}[\mathbf{y}|\mathbf{x}]$, and suppose for now that $p$ satisfies $(\mathcal{C},0)$-multiaccuracy for $\mathcal{D}$. Let $\tilde{p}^{*}=2p^{*}-1$ and $\tilde{p}=2p-1$ denote the [1,1][-1,1] versions of the predictors $p^{*}$ and $p$. Let $\mathsf{span}(\mathcal{C})$ denote the vector space spanned by $\mathcal{C}$. We can write,
@@
where $\tilde{p}_{C},\tilde{p}^{*}_{C}$ are the components of $\tilde{p}$ and $\tilde{p}^{*}$ in the $\mathsf{span}(\mathcal{C})$ respectively, and $\tilde{p}_{\perp},\tilde{p}^{*}_{\perp}$ are the components of $\tilde{p},\tilde{p}^{*}$ orthogonal to $\mathsf{span}(\mathcal{C})$. We recall that the inner product $\langle f,g\rangle$ for functions $f,g:\mathcal{X}\to\mathbb{R}$ is defined by $\operatorname*{\mathbf{E}}_{\mathcal{D}_{\mathcal{X}}}[f(\mathbf{x})g(\mathbf{x})]$. Since $p$ is $(\mathcal{C},0)$-multiaccurate for $\mathcal{D}$, we have that,
$\langle c,\tilde{p}-\tilde{p}^{*}\rangle=2\operatorname*{\mathbf{E}}[c(\mathbf{x})(p(\mathbf{x})-p^{*}(\mathbf{x}))]=0$
@@
As $\tilde{p}_{\perp},\tilde{p}^{*}_{\perp}$ are both orthogonal to $\mathsf{span}(\mathcal{C})$, and $\tilde{p}_{C}-\tilde{p}^{*}_{C}\in\mathsf{span}(\mathcal{C})$, we must have $\tilde{p}_{C}=\tilde{p}^{*}_{C}$. What this suggests is that even though technically we can have
@@
$\mathrm{cor}(\mathbf{y},\tilde{p}_{C})=\langle\tilde{p}^{*},\tilde{p}_{C}\rangle$ $=\langle\tilde{p}^{*}_{C}+\tilde{p}^{*}_{\perp},\tilde{p}_{C}\rangle$
@@
Above in the last line we used that $\tilde{p}^{*}_{\perp}$ is orthogonal to $\mathsf{span}(\mathcal{C})$ and that $\tilde{p}_{C}\in\mathsf{span}(\mathcal{C})$, as well as the observation that $\tilde{p}_{C}=\tilde{p}^{*}_{C}$. Thus, a projection of $\tilde{p}$ onto $\mathsf{span}(\mathcal{C})$ could potentially lead to a weak agnostic learner. However, we need to take some care as (a) $\tilde{p}_{C}$ need not have range [1,1][-1,1] (or be bounded even), and (b) if we don’t have $(\mathcal{C},0)$-multiaccuracy, but only $(\mathcal{C},\tau)$, then depending on the representation of $\tilde{p}_{C}$ in terms of $c\in C$, the errors could accumulate making the above analysis trivial. Observe that this constitutes a much richer form of post-processing that considered in Section 4.
@@
For simplicity, let $\tilde{p}=2p-1$ denote the [1,1][-1,1]-version of $p$. Then,
@@
Above in the second last inequality, we used the fact that $p$ is $(\mathcal{C},\tau)$-multiaccurate for $\mathcal{D}$ and that $c_{s}\in\mathcal{C}$ for all $s$, and in the last inequality we use that clipping $q$ to get $h\in[-1,1]$ only reduces the squared error with respect to $2\mathbf{y}-1$ and that $2\mathbf{y}-1\in\{-1,1\}$. The conclusion then follows by rearranging. ∎
@@
for $\gamma\geq 4\tau\sum_{s}|\lambda_{s}|+\beta$. Then, $h=\operatorname{clip}(q)$, which clips $q$ to the range [1,1][-1,1], satisfies $\mathrm{cor}_{\mathcal{D}}(\mathbf{y},h)\geq\beta$.
@@
In the last inequality, we used $\mathrm{cor}(p,c)\geq\alpha$ and $\|c\|\leq 1$, as $c$ has range [1,1][-1,1].